OpenStudy (anonymous):

If I am given the relationship between g(x) and f(x), how do I find the relationship between g'(x) and f'(x)? I am given g(x) = f(3x) What is the relationship between g'(x) and f'(x)?

jimthompson5910 (jim_thompson5910):

well if g(x) = f(x), you agree that g ' (x) = f ' (x) right?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

the only thing that's different is the 'x' in f(x) is 3x

jimthompson5910 (jim_thompson5910):

so we use the chain rule g(x) = f(3x) g ' (x) = f ' (3x)*d/dx[3x] g ' (x) = f ' (3x)*3 g ' (x) = 3 * f ' (3x)

jimthompson5910 (jim_thompson5910):

the notation d/dx[3x] means "take the derivative of 3x with respect to x"

OpenStudy (anonymous):

Oh yea, that make a lot of sense. Thank you.

jimthompson5910 (jim_thompson5910):

no problem